Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis ofComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Req 3A Req 3B Reg 4 Prepare a standard cost caComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Req 3A Req 3B Reg 4 Compute the standard direcComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req 3A Reg 3B Req 4 Complete the following ManReq 1 Reg 2 Req ЗА Req 3B Rita 4 Determine the reason for any underapplied or overapplied overhead for the year by com and ef
Answer #1

Required 1:-

Predetermined overhead rate $14.00 per DLH
Variable rate $5.00 per DLH
Fixed rate $9.00 per DLH

Working:-

Fixed rate:- budgeted fixed overhead/Estimated direct labour hour

= 2295000/255000 = 9.00

Predetermined overhead rate = 5.00+9.00 = 14.00

Required 2:-

Direct material 4 pounds at 10.50 per pound $42.00
Direct labour 1.5 DLHs at 13.50 per DLH $20.25
VaroaVar overhead 1.5 DLHs at 5.00 per DLH $7.50
Fixed overhead 1.5 DLHs at 9.00 per DLH $13.50

Total = $83.25

Required 3a:-

Standard direct labour hour required for actual production = 204000units * 1.5 hrs. = 306000 units

Required 3b :-

Manufacturing overhead

Actual cost 3613350 WIP (306000*14.00) 4284000
Overapplied oberhove 670650

Required 4:-

Variable overhead variance $696150 F
Variable overhead efficiency variance $127500 U
Fixed overhead budgeted variance $357000 U
Fixed overhead volume variance $459000 F

Working:-

a. Variable overhead variance :- (AR-SR)*AH

= (5.00-2.9)*331500 = 696150

b. Variable overhead efficiency variance = (SH-AH)*SR

= (306000-331500)*5.00 = 127500

c. Fixed overhead budgeted variance =

Budgeted overhead - Actual overhead

2295000-2652000 = 357000

d.Fixed overhead volume variance =

Overhead applied - budgeted overhead

(306000*9.00)- 2295000 = 459000

SR standard rate
AR actual rate
SH standard hour
AH actual hour